### v1

# ReZero: Boosting MCTS-based Algorithms by Just-in-Time and Speedy Reanalyze

Abstract
MCTS-based algorithms, such as MuZero and its derivatives, have achieved widespread success in various decision-making domains. These algorithms employ the reanalyze process to enhance sample efficiency, albeit at the expense of significant wall-clock time consumption. To address this issue, we propose a general approach named ReZero to boost MCTS-based algorithms. Specifically, we propose a new scheme that simplifies data collecting and reanalyzing, which significantly reduces the search cost while guarantees the performance as well. Furthermore, to accelerate each search process, we conceive a method to reuse the subsequent information in the trajectory. The corresponding analysis conducted on the bandit model also provides auxiliary theoretical substantiation for our design. Experiments conducted on Atari environments and board games demonstrates that ReZero substantially improves training speed while maintaining high sample efficiency. The code is available as part of the LightZero benchmark at https://github.com/opendilab/LightZero.
1 Introduction
As a pivotal subset of artificial intelligence, Reinforcement Learning (RL) (Sutton & Barto, 1988) has acquired achievements and applications across diverse fields, including interactive gaming (Vinyals et al., 2019), autonomous vehicles (Li et al., 2022), and natural language processing (Rafailov et al., 2023). Despite its successes, a fundamental challenge plaguing traditional model-free RL algorithms is their low sample efficiency. These algorithms typically require a large amount of data to learn effectively, making them often infeasible in the real world. In response to this problem, numerous model-based RL(Janner et al., 2019) (Hafner et al., 2020) methods have emerged. These methods involve learning an additional model of the environment from data and utilizing the model to assist the agent’s learning, thereby improving sample efficiency significantly.
Within this field, Monte Carlo Tree Search (Świechowski et al., 2023) (MCTS) has been proven to be a efficient method for utilizing models for planning. It incorporates the UCB1 algorithm (Auer et al., 2002) into the tree search process and has achieved promising results in in a wide range of application scenarios. Specifically, AlphaZero (Silver et al., 2017) plays a big role in combining deep RL with MCTS, achieving notable accomplishments that can beat top-level human players. While AlphaZero can only be applied to environments with perfect simulators, MuZero (Schrittwieser et al., 2019) extended the algorithm to cases without known environment models, resulting in good performances in a wider range of tasks. Following MuZero, many successor algorithms have emerged, enabling MuZero to be applied in continuous action spaces (Hubert et al., 2021), environments with stochastic dynamics (Antonoglou et al., 2021), offline RL training scenarios (Schrittwieser et al., 2021), and etc. All these MCTS-based algorithms (Silver et al., 2016, 2017; Schrittwieser et al., 2019; Ye et al., 2021) made valuable contributions to the universal applicability of the MCTS+RL framework.
However, the extensive tree search computations incur additional time overhead for these algorithms: During the data collection phase, the agent needs to execute MCTS to select an action every time it receives a new state. Furthermore, due to the characteristics of tree search, it is challenging to parallelize it using commonly used vectorized environments methods, further amplifying the speed disadvantage. On the other hand, during the reanalyze (Schrittwieser et al., 2021) process, in order to obtain higher-quality update targets, the latest model parameters are used to re-run MCTS on samples in buffer. The time required to process samples thus increases as a trade-off for high sample efficiency. The excessive time cost has become a bottleneck hindering the further promotion of these algorithms.
Recently, a segment of research endeavors are directed toward mitigating the above wall-clock time overhead about MCTS-based algorithms. On the one hand, SpeedyZero (Mei et al., 2023) greatly diminishes algorithms’ time overheads by deploying a parallel system; however, it demands more computational resources. On the other hand, it remains imperative to identify methodologies that accelerate these algorithms without imposing extra computational demands. PTSAZero (Fu et al., 2023), for instance, compresses the search space via state abstraction, decreasing the time cost per search. Nevertheless, it is widely acknowledged that MCTS-based algorithms necessitate search times to assure their performance. From this perspective, our research try to discover strategies that preserve algorithmic performance while minimizing the search times as much as possible. Moreover, given the observation that the tree search trajectories for states at adjacent time-steps frequently exhibit a significant degree of overlap, it becomes feasible to leverage the reuse of search information across different time-steps as a method to further expedite each search process.
In this paper, we introduce a new approach designed to boost the MCTS-based algorithms. Firstly, drawing inspiration from the utilization of a periodically updated target network in DQN (Mnih et al., 2013), we find that it is not essential to reanalyze the mini-batch at each iteration. Therefore, we propose a simplified framework which periodically reanalyzes the buffer to update policy targets in a just-in-time manner. Such an approach markedly diminishes the number of calls to MCTS. Secondly, within the reanalyze process, given the collected game trajectory, we propose to reuse search information from subsequent state to accelerate the search of the current state. In addition, we provide theoretical analysis on the potential implications of these modifications about the quality of target policy. Our approach can be easily applied to the MCTS-based algorithm family and can be synergistically implemented alongside existing acceleration methods. In addition, it will not bring about an increase in hardware overhead. Empirical experiments demonstrate that our approach yields good results in both single-agent environment (Atari (Bellemare et al., 2013)) and two-player board games (Silver et al., 2016), greatly improving the training speed while maintaining high sample efficiency. Ablation experiments explore the impact of different reanalyze periods and the acceleration effect of information reuse on a single search. The main contributions of this paper can be summarized as follows:
  • We propose a simplified framework that greatly reduces the number of calls to MCTS, thus speeding up MCTS-based algorithms.
  • We propose a method to speed up a single search by reusing search information. Theoretical support for the effectiveness of the method is also provided.
  • We conduct serveal experiments on diverse environments and further investigate our method through ablations.
2 Related Work
2.1 MCTS-based Algorithms
AlphaGo (Silver et al., 2016) and AlphaZero (Silver et al., 2017) combined MCTS with deep reinforcement learning, achieving significant results in board games, repeatedly defeating world-champion players, and revealing their remarkable capabilities. Following these achievements, MuZero (Schrittwieser et al., 2019) algorithm combines tree search with a learned value equivalent model (Grimm et al., 2020b), successfully promoting the algorithm’s application to scenarios without known models, such as Atari. Subsequent to MuZero’s development, several new works based on the MuZero framework emerged. Ye et al. (2021) further improved MuZero’s sample efficiency, using very little data for training and still achieving outstanding performance. Hubert et al. (2021) and Antonoglou et al. (2021) extended the MuZero agent to environments with large action spaces and stochastic environments, further broadening the algorithm’s usage scenario. Schrittwieser et al. (2021) applied MuZero’s reanalyze operation as a policy enhancement operator and extended it to offline settings. Danihelka et al. (2022) changed the node selection method during tree search, so the policy obtained from the search shows a monotonic improvement, ensuring algorithm performance even with low simulation numbers. Despite the aforementioned excellent improvement techniques, the current time overhead of MCTS-based RL remains significant, which is the main problem this paper aims to address.
2.2 MCTS Acceleration
Recent research has focused on accelerating MCTS-based algorithms. The work by Mei et al. (2023) significantly reduces the algorithm’s time overhead by designing a parallel system; however, this approach requires more computational resources and involves some adjustments for large batch training. Fu et al. (2023) narrows the search space through state abstraction, which amalgamates redundant spatial information, thereby reducing the time cost per search. Silver et al. (2016); Wu (2019) use information reuse to speed up the search process. They save sub-trees of the search tree and serve as initialization for the next search. However, this naive front-to-back information reuse is significantly different from our approach. To our knowledge, we are the first to enhance MCTS by reusing information in a backwards manner. Our proposed approach can seamlessly integrate with various MCTS-based algorithms, many of which (Danihelka et al., 2022; Mei et al., 2023; Fu et al., 2023; Leurent & Maillard, 2020) are orthogonal to our contributions.
3 Preliminaries
3.1 MuZero
MuZero (Schrittwieser et al., 2019) is a fundamental model-based RL algorithm that incorporates a value-equivalent model (Grimm et al., 2020a) and leverages Monte Carlo Tree Search for planning within a learned latent space. The model consists of three core components: a representation model $h_{\theta}$, a dynamics model $g_{\theta}$, and a prediction model $f_{\theta}$:
$\begin{array}[]{lll}\text{Representation:}&&s_{t}=h_{\theta}(o_{t-l:t})\\ \text{Dynamics:}&&s_{t+1},r_{t}=g_{\theta}(s_{t},a_{t})\\ \text{Prediction: }&&v_{t},p_{t}=f_{\theta}(s_{t})\\ \end{array}$ (1)
During the inference phase, the representation model transforms a sequence of the last $l$ observations $o_{t-l:t}$ into a corresponding latent state representation $s_{t}$. The dynamics model processes this latent state alongside an action $a_{t}$, yielding the subsequent latent state $s_{t+1}$ and an estimated reward $r_{t}$. Finally, the prediction model accepts a latent state and produces both the predicted policy $p_{t}$ and the state’s value estimate $v_{t}$. These outputs are instrumental in guiding the agent’s action selection process throughout its MCTS. Lastly the agent selects or samples the best action $a_{t}$ following the searched visit count distribution. During the training phase, given a training sequence $\{o_{t-l:t+K},a_{t+1:t+K},u_{t+1:t+K},\pi_{t+1:t+K},z_{t+1:t+K}\}$ at time $t$ sampled from the replay buffer, where $u_{t+k}$ denotes the actual reward obtained from the environment, $\pi_{t+k}$ represents the target policy obtained through MCTS during the agent-environment interaction, and $z_{t+k}$ is the value target computed using n-step bootstrapping (Hessel et al., 2018). The representation model initially converts the sequence of observations $o_{t-l:t}$ into the latent state $s_{t}^{0}$. Subsequently, the dynamic model executes $K$ latent space rollouts based on the sequence of actions $a_{t+1:t+K}$. The latent state derived after the $k$-th rollout is denoted as $s_{t}^{k}$, with the corresponding predicted reward indicated as $r_{t}^{k}$. Upon receiving $s_{t}^{k}$, the prediction model generates a predicted policy $p_{t}^{k}$ and a estimated value $v_{t}^{k}$. The final training loss encompasses three components: the policy loss ($l_{p}$), the value loss ($l_{v}$), and the reward loss ($l_{r}$):
$L_{\text{MuZero}}=$ $\sum_{k=0}^{K}l_{p}(\pi_{t+k},p_{t}^{k})+\sum_{k=0}^{K}l_{v}(z_{t+k},v_{t}^{k})+\sum_{k=1}^{K}l_{r}(u_{t+k},r_{t}^{k})$ (2)
MuZero Reanalyze, as introduced in (Schrittwieser et al., 2021), is an advanced iteration of the original MuZero algorithm. This variant enhances the model’s accuracy by conducting a fresh Monte Carlo Tree Search on sampled states with the most recent version of the model, subsequently utilizing the refined policy from this search to update the policy targets. Such reanalysis yields targets of superior quality compared to those obtained during the initial data collection phase. The Schrittwieser et al. (2021) expands upon this approach, formalizing it as a standalone method for policy refinement. This innovation opens avenues for its application in offline settings, where interactions with the environment are not possible. Traditional uses of the algorithm have intertwined reanalysis with training, periodically revisiting a subset of samples within a mini-batch for reevaluation. Furthermore, we suggests a paradigm shift, advocating for the decoupling of the reanalyze process from training routines, thus providing a more flexible and potentially more efficient methodology.
3.2 Stochastic Bandits
In the upcoming Section 4.2.2, we plan to offer an in-depth look at the theoretical assurances behind the performance of our newly designed ReZero agent, which is rooted in the fundamentals of bandit theory. Before diving into that, we’ll start with an easy-to-understand overview of the key concepts surrounding Stochastic Bandits.
A stochastic bandit $\nu$ with $K$ arms (Simchi-Levi et al., 2023) can be formulated as $\nu=\{P_{i}|i\in{1,2,...,K}\}$, where $P_{i}$ denotes the reward distribution corresponding to the arm $i$. The expected reward of arm $i$ is denoted as $\mu_{i}(\nu)$. Given an algorithm $\pi$ interacting with the bandit $\nu$ for $n$ rounds, we define the regret $R_{n}$ as follows:
$R_{n}(\pi,\nu)=n\mu^{*}(\nu)-\mathbb{E}[\sum_{t=1}^{n}X_{t}]$ (3)
where $\mu^{*}(\nu)=\max_{a\in[K]}\mu_{a}(\nu)$, $X_{t}$ denotes the random reward achieved at time $t$.
The regret quantifies the loss of reward incurred by $\pi$ compared to the optimal policy on $\nu$ during $n$ rounds of interactions. In practical problems, we aim for $\pi$ to exhibit sub-linear regret growth across a category of bandits $\Omega$ , i.e.:
$\lim_{n\to\infty}\frac{R_{n}(\pi,\nu)}{n}=0,\ \ \ \text{for all $\nu\in\Omega$ }$ (4)
where sub-linear growth rate indicates $\pi$ is choosing the optimal arm almost all of the time with $n$ tends to infinity. UCB (Auer et al., 2002) is a classical bandit algorithm that achieves regret growth rate of $O(\log(n))$.
AlphaZero (Silver et al., 2017) considers the child nodes in tree search as bandit arms and applies a variant of the UCB in the tree search process:
$UCB_{score}(s,a)=Q(s,a)+cP(s,a)\frac{\sqrt{{\sum}_{b}N(s,b)}}{1+N(s,a)}$ (5)
where $s$ is the state corresponding to the current node, $a$ is the action corresponding to a child node. $Q(s,a)$ is the mean return of choosing action $a$, $P(s,a)$ is the prior score of action $a$, $N(s,a)$ is the total time that $a$ has been chosen, ${\sum}_{b}N(s,b)$ is the total time that $s$ has been visited.
4 Method
Figure 1: Execution workflow and runtime cycle graph about MuZero and ReZero in both single and multiple worker cases. The number inside the modules represent the number of iterations, and the number under the modules represent the time required for module execution. The model is updated $n$ iterations between two collections. MuZero reanalyzes the mini-batch before each model update. ReZero reanalyzes the entire buffer after certain update iterations($\frac{n}{3}$ for example), which not only reduces the total number of MCTS calls, but also takes advantage of the processing speed of large batches.
In this section, we introduce two techniques to boost MCTS-based algorithms. Section 4.1 describes the design of just-in-time reanalyze, which achieves sufficient training on policy target meanwhile significantly reduces the number of MCTS executions. Furthermore, section 4.2 introduces speedy reanalyze: an acceleration method based on information reuse, which utilizes search results from the subsequent time-step to speed up the search at the previous time-step, thereby saving the time required for a single MCTS execution.
4.1 Just-in-Time Reanalyze: A Simplified Framework
Reanalyze, as a method to enhance the quality of training targets, was previously bound together with the training process in earlier implementations (Schrittwieser et al., 2021) (Figure 1). However, recomputing the target policy at each training iteration incurs massive MCTS computations. This dilemma motivates us to find a method to reduce the amount of MCTS calculations while ensuring performance.
We begin our detailed analysis with the following perspective: consider the reanalyze process as a policy improvement operator $\mathcal{R}_{\theta^{t}}$, which takes the current policy $P_{\theta^{t}}$ and transforms it into the target policy $T_{\theta^{t}}$ for training.
$T_{\theta^{t}}(a|s)=(\mathcal{R}_{\theta^{t}}P_{\theta^{t}})(a|s)=\frac{E_{T_{a}\sim\theta^{t}}[T_{a}]}{N}$ (6)
where $N$ is the number of simulations during the tree search and $E[T_{a}]$ is the expected total number that action $a$ was chosen. During the training process, we update the model parameter $\theta$ according to the following equation:
$\theta^{t+1}=\arg\min_{\theta}E_{s\sim\mathcal{D}}[\sum_{k=0}^{K}l^{p}(T_{\theta^{t}}(\cdot|s^{k})),P_{\theta}(\cdot|s^{k}))+l_{reg}]$ (7)
where $s$ is the state sampled from buffer $\mathcal{D}$, and $s^{k}$ is the state after $k$ unroll steps. We consider $l_{r}$ and $l_{v}$ in Equation 2 as a regularizer for $l_{p}$, and mark them as $l_{reg}$. In this way, the training can be seen as a projection operator $\mathcal{P}$ (Grill et al., 2020), projecting $T_{\theta^{t}}$ onto the policy space. The alternation of the reanalyze process and the training process constitutes policy iteration (Sutton & Barto, 1988), which can be seen as the composition of two operators:
$P_{\theta^{t+1}}=(\mathcal{P}\circ\mathcal{R}_{\theta^{t}})P_{\theta^{t}}$ (8)
In previous work (Schrittwieser et al., 2021), the reanalyze process is executed on the mini-batch of each training iteration. Changing the projection target frequently may result in an insufficient optimization, thereby degrading the quality of $\mathcal{P}$. In this way, even if the improvement operator $\mathcal{R}_{\theta}$ can produce high-quality training targets, the composite operator may not perform well. Therefore, we can adopt the practice of periodically updating the target network in DQN (Mnih et al., 2013), using Just-in-Time Reanalyze to change the projection targets only after several rounds of training.
Specifically, we propose the periodical reanalyze mechanism, i.e., we reanalyze the whole buffer after a fixed number of iterations. For each training iteration, we only need to sample the mini-batch and execute the gradient descent. Compared with previous methods, our method greatly reduces the number of MCTS, and can improve the parallelism due to the bigger search batch (Keskar et al., 2016), further enhanced the overall speed. Since the reanalyze process provides the target policy, the tree search during the collection phase is no longer necessary in the entire pipeline. Thus, we can bypass the tree search and directly select actions based on the policy, further reducing MCTS executions.
Based on our simplified framework, there are several potential advantages that may be explored in future work:
  • When directly using policy for data collection, action selection is no longer constrained by the tree search. As a result, previous vectorized environments such as (Weng et al., 2022) can be integrated to speed up. Besides, this design also makes MCTS-based algorithms compatible to existing RL exploration methods like (Badia et al., 2020).
  • Our method no longer needs to reanalyze the mini-batch for each iteration, thus decoupling the process of reanalyze and training. This provides greater scope for parallelization. In the case of multiple workers, we can design efficient parallelization paradigms as shown in Figure 1.
4.2 Speedy Reanalyze: Temporal Information Reuse
The above framework reduces the total number of MCTS. Next, we propose an temporal information reuse method to speed up the single tree search (Hereafter, termed information reuse for brevity). Section 4.2.1 introduces motivations and specific details. Section 4.2.2 carries out theoretical analysis of the method based on the bandit model.
4.2.1 Reuse the Information from Backward
As a local planning approach (Świechowski et al., 2023), MCTS is executed when receives each environment’s state. The agent, by performing tree search on the current state, can gain the root value and the visit counts: the former tells the quality of the state, and the later tells which actions are more likely to yield good returns. These can be seen as a kind of information. The trajectories traversed by the search of adjacent time-steps often have many overlaps, indicating that it involves similar knowledge. If the information gained at one time-step can be passed on to another, the search process can be accelerated: for example, if the search conducted at time-step $A$ reveals the quality of state $s$, then at time-step $B$, the agent need not to spend extra search cost to explore state $s$. Therefore, we designed a method to speed up the reanalyze process by reusing search information.
Figure 2: An illustration about the information reuse between adjacent time-steps. We search in the reverse order of the trajectory. After performing search on $s_{t+1}$, we obtain the search tree $T_{t+1}$ and the root value $m_{t+1}$. When searching on $s_{t}$, we set the value of the node corresponding to $s_{t+1}$ in search tree $T_{t}$ as $m_{t+1}$. If action $a_{t}$ is selected, the traverse process will be terminated, otherwise the traverse will proceed normally.
Algorithm 1 shows the specific design with Python-like code. During reanalyze, since the trajectories was already collected in the replay buffer, we can perform the search in a reverse order. That is to say, we first perform tree search on the state of the last time-step in the trajectory, and then search the second to last, and so on. After we search on state $s_{t+1}$  [^1] To facilitate the explanation, we have omitted details such as the latent space and do not make a deliberate distinction between states and nodes. (the state of the $t+1$-th time-step in the trajectory), the root value $m_{t+1}$ is obtained. When performing search on $s_{t}$, we set the value of $s_{t+1}$ to the fixed value $m_{t+1}$. During traverse in the tree, we select the action $a_{root}$ for root node $s_{t}$ according to the following equation:
$a_{root}=\arg\max_{a}I_{t}(a)$ (9)
$I_{t}(a)=\begin{cases}UCB_{score}(s_{t},a),&a\neq a_{t}\\ r_{t}+\gamma m_{t+1},&a=a_{t}\end{cases}$ (10)
where $a$ is an action corresponding to a child node, $a_{t}$ is the action corresponding to $s_{t+1}$, and $r_{t}$ is the reward predicted by the dynamic model. If an action other than $a_{t}$ is selected, the simulation just traverses following the original setting in MCTS (Schrittwieser et al., 2019). If action $a_{t}$ is selected, this simulation is terminated immediately. Since the time used to search for node $s_{t+1}$ is saved, the modified search process is faster than the original version, achieving Speedy Reanalyze.
Algorithm 1 Python-like code for information reuse
# trajectory_segment: a segment with length K
def search_backwards(trajectory_segment):
# prepare search context from the segment
roots, actions = prepare(trajectory_segment)
policy_targets = []
# search the roots backwards
for i in range(K, 1, -1):
if i == K:
# origin MCTS for Kth root
policy, value = origin_MCTS(roots[i])
else:
# reuse information from previous search
policy, value = reuse_MCTS(
roots[i], actions[i], value
policy_targets.append(policy)
# N: simulation numbers during one search
def reuse_MCTS(root, action, values)
for i in range(N):
# select an action for root node
a = select_root_child(root, action, value):
# early stop the simulation
if a == action:
backpropagate()
break
# traverse to the leaf node
else:
traverse()
backpropagate()
4.2.2 Theoretical Analysis
AlphaZero employs a variant of the UCB algorithm during MCTS search. Subsequently, MuZero proposes further improvements to this variant formula. Practical evidence has demonstrated the effectiveness of both these node selection methods; however, they have not been theoretically analyzed in their original paper. In earlier studies, UCT analyzed the convergence properties of the UCB1 algorithm in a tree structure, albeit without introducing a prior policy. PUCB (Rosin, 2011) incorporates relevant analysis of a prior policy, but it uses the prior in a penalty function-like approach which is different from in the MCTS-based algorithm family. In recent work, Gumbel MuZero (Danihelka et al., 2022) approximated the tree search model as a stochastic bandit, and on this basis, proved the monotonic improvement of their designs. In this section, we will draw upon such a simplification and demonstrate within the context of a stochastic bandit model why the node selection formula and policy updating methods used in the MCTS-based algorithm family are effective, and how information reuse can influence the policy target.
Definition 4.1.
After $n$ simulations with a stochastic bandit $\nu$ with $K$ arms, let $T_{i}$ denote the number of visits to arm $i$, then the policy generated based on the visit distribution is a random vector $P=[\frac{T_{1}}{n},...,\frac{T_{K}}{n}]$ The value $V_{\nu}(P)$ of policy $P$ on $\nu$ is defined as:
$V_{\nu}(P)=\sum_{i}\frac{T_{i}\mu_{i}(\nu)}{n}$ (11)
In cases where it does not cause any ambiguity, we will omit the mark $\nu$ in the following text. The next theorem explains the rationale behind using distribution as a target policy:
Theorem 4.2.
If an algorithm $\pi$ exhibits sub-linear growth in regret $R_{n}$ with respect to a bandit $\nu$, then the expectation $E[V(P)]$ converges to $\mu^{*}$ as $n$ tends to infinity.
$E[V(P)]$ $=\sum_{i}\frac{E[T_{i}]\mu_{i}}{n}$ (12)
$=\sum_{i}\frac{E[T_{i}](\mu^{*}-\Delta_{i})}{n}$ (13)
$=\mu^{*}-\frac{\sum_{i}E[T_{i}]\Delta_{i}}{n}$ (14)
where $\Delta_{i}=\mu^{*}-\mu_{i}$, and $\sum_{i}E[T_{i}]\Delta_{i}$ is the regret $R_{n}$ of the algorithm (see Appendix A).
Since $R_{n}$ growths in a sub-linear rate, so we can get
$\lim_{n\to\infty}E[V(P)]=\mu^{*}-\lim_{n\to\infty}\frac{R_{n}}{n}=\mu^{*}$ (15)
The following theorem provides an upper bound for the growth rate of $R_{n}$ for choosing arm in AlphaZero’s way , thereby demonstrating that the regret is sub-linearly increasing.
Theorem 4.3.
Denote the prior for arm $i$ as $P_{i}$. Without loss of generality, suppose $\mu^{*}=\mu_{1}$. Assume that the reward distribution of each arm $i$ is $1$-subgaussian and $P_{1}>\mu_{1}$, then the regret $R_{n}$ of AlphaZero satisfies (Appendix A)
$R_{n}\leq\sum_{i:\Delta_{i}>0}\inf_{\varepsilon\in(0,\Delta_{i})}\Delta_{i}(3+\frac{16}{\varepsilon^{4}P_{1}^{2}}+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon})$ (16)
The upper bound given by Theorem 4.3 is $O(\sqrt{n})$, which, while not reaching the level of $O(\log{n})$, is still sub-linear. Inequality (16) also indicates that a better policy which has higher probability to choose the best arm can lower the upper bound. Combining 4.2 and 4.3, we can conclude that with the increase in the number of simulations, the visit distribution will increasingly concentrate on the optimal arm, and the value corresponding to the policy will also converge to the value of the optimal arm. If the assumption that $P_{1}>\mu_{1}$ doesn’t hold, an upper bound with $O(\sqrt{n})$ rate can still be obtained by similar proof as long as there exists a $t$ that $P_{1}\sqrt{t-1}>\mu_{1}$, which indicates unlike some algorithms that are sensitive to prior policies, MCTS has strong robustness to prior policies. As long as the number of simulations is sufficient, it can still achieve a good target policy.
When reusing information in the search process, we can obtain a reference value for an arm. The existence of this reference value can eliminate the need for exploration of known arms, thereby further reducing the algorithm’s regret. The following theorem proves this point.
Theorem 4.4.
For a bandit which satisfies the assumptions in 4.3 and an already known arm value, the regret upper bound can be reduced if we use the way of 4.2.1 to choose arm. Specifically, if the known arm is the optimal arm with value $\mu_{1}$, then the regret $R_{n}$ satisfies:
$R_{n}\leq\sum_{i:\Delta_{i}>0}\inf_{\varepsilon\in(0,\Delta_{i})}\Delta_{i}(2+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon})$ (17)
In other case, with the known value $\mu_{l}$ of a sub-optimal arm $l$, the regret $R_{n}$ satisfies:
$R_{n}$ $\leq\sum_{i:\Delta_{i}>0,i\neq l}\inf_{\varepsilon\in(0,\Delta_{i})}\Delta_{i}(3+\frac{16}{\varepsilon^{4}P_{1}^{2}}$ (18)
$+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon})$ (19)
$+\inf_{\varepsilon\in(0,\Delta_{l})}\Delta_{l}(1+\frac{16}{\varepsilon^{4}P_{1}^{2}})$ (20)
The complete proof can be found in Appendix A.
Figure 3: Time-efficiency of ReZero-M vs. MuZero on six representative Atari games and two board games (Connect4, Gomoku). The horizontal axis represents Wall-clock Time (hours), while the vertical axis indicates the Episode Return over 5 episodes. ReZero-M achieves higher time-efficiency than the baseline on most Atari games and Gomoku, and achieve similar time-efficiency on some simpler game like Connect4. Mean of 5 runs, shaded areas are 95% confidence intervals.
Atari Board Games
avg. wall time (h) to 100k env. steps $\downarrow$ Pong Breakout MsPacman Qbert Seaquest UpNDown Connet4 Gomoku
ReZero-M (ours) $\mathbf{1}\scriptstyle{\pm 0.6}$ $\mathbf{7.7}\scriptstyle{\pm 0.3}$ $\mathbf{3.8}\scriptstyle{\pm 0.2}$ $\mathbf{5.1\scriptstyle{\pm 0.3}}$ $\mathbf{6.1}\scriptstyle{\pm 0.5}$ $\mathbf{6.9}\scriptstyle{\pm 0.5}$ $\mathbf{10}\scriptstyle{\pm 0.8}$ $\mathbf{9\scriptstyle{\pm 0.8}}$
MuZero (Schrittwieser et al., 2019) $4.1\scriptstyle{\pm 0.2}$ $12.4\scriptstyle{\pm 0.8}$ $12.0\scriptstyle{\pm 0.9}$ $10.2\scriptstyle{\pm 0.5}$ $11.75\scriptstyle{\pm 0.9}$ $12.4\scriptstyle{\pm 0.9}$ $11\scriptstyle{\pm 0.9}$ $23\scriptstyle{\pm 2}$
Table 1: Average wall-time of ReZero-M vs. MuZero on various tasks. (left) Six Atari games, (right) two board games. The time represents the average total wall-time to 100k environment steps for each algorithm. Mean and standard deviation over 5 runs. ReZero used two to four times less wall-time per 100k steps than the baseline method.
From the proof, it can be seen that if the real value $\mu_{1}$ of the optimal arm is known in advance, the regret caused by the underestimation of the optimal arm can be eliminated. However, the part of the regret that is reduced is $O(1)$. If the real value $\mu_{i}$ of a sub-optimal arm $l$ is known in advance, the regret caused by continuous exploration of this sub-optimal arm can be eliminated. The part of the reduced regret is $O(\sqrt{n})$. Intuitively, it can be understood this way: If the value of a sub-optimal arm $l$ is known, then after sufficient exploration of the optimal arm, the upper confidence bound will converge to be higher than $\mu_{l}$ with probability $1$, thus making the regret caused by arm $l$ no longer increases with $n$. While if the value of the optimal arm is known, then as the number of simulations $n$ increases, there will always be a moment when the upper confidence bound of the sub-optimal arm $i$ is greater than $\mu_{1}$, so arm $i$ will still be repeatedly selected, leading to the regret caused by arm $i$ increases with $n$. However, in this case, since the optimal arm is chosen the most time as $n$ increases, we can save more time by early stopping the search.
5 Experiment
Efficiency in RL usually refers to two aspects: sample efficiency, the agent’s ability to learn effectively from a limited number of environmental interactions, gauged by the samples needed to reach a successful policy using equal compute resources; time efficiency, which is how swiftly an algorithm learns to make optimal decisions, indicated by the wall-clock time taken to achieve a successful policy. Our primary goal is to improve time efficiency without compromising sample efficiency. To empirically validate the efficiency of the two boosting techniques introduced in Section 3.2 for MCTS-based algorithms, we have incorporated the ReZero with two prominent algorithms detailed in (Niu et al., 2023): MuZero with a Self-Supervised Learning loss (SSL) and EfficientZero, yielding the enhanced variants ReZero-M and ReZero-E respectively. For brevity, when the context is unambiguous, we simply refer to them as ReZero, omitting the symbol representing the baseline algorithm.
In order to validate the applicability of our method across multiple types of decision-making environments, we opted for six representative Atari environments characterized by classic image-input observation and discrete action spaces, in addition to the strategic board games Connect4 and Gomoku with special state spaces. For a baseline comparison, we employed the original implementations of MuZero and EfficientZero as delineated in the LightZero benchmark (Niu et al., 2023). The full implementation details available in Appendix B. Besides, we emphasize that to ensure an fair comparison of wall-clock time, all experimental trials were executed on a fixed single worker hardware settings.
In the next subsections, we sequentially explore and try to answer the following three questions:
  • How much can ReZero-M/ReZero-E improve time efficiency compared to MuZero/EfficientZero, while maintaining equivalent levels of high sample efficiency?
  • What is the concrete effect and hyper-parameter sensitivity of the Just-in-Time Reanalyze technique in ReZero?
  • How much search budgets and time costs can be saved in practice by the Temporal Information Reuse technique?
5.1 Time Efficiency
Setup In this section, we aim to evaluate the performance of ReZero against classical MCTS-based algorithms MuZero and EfficientZero, focusing on the wall-clock time reduction required to achieve the comparable performance level. In order to facilitate a fair comparison in terms of wall-time, for each run, we not only utilize identical computational resources but also maintain consistent hyper-parameter settings across the algorithms (unless otherwise specified cases). Key parameters are aligned with those from original papers. Specifically, we set the replay ratio (the ratio between environment steps and training steps) (D’Oro et al., 2022; Schwarzer et al., 2023) to 0.25, and the reanalyze ratio (the ratio between targets computed from interactions with the environment and by reanalysing existing data) (Schrittwieser et al., 2021) is set to 1. For detailed hyper-paramater configurations, please refer to the Appendix B.
Results Figure 3 illustrates the performance comparison in terms of wall-clock time between ReZero-M and the original MuZero across a variety of RL environments. It is apparent that ReZero-M achieves a significant improvement in time efficiency, reaching a near-optimal policy in substantially less time for the games Pong, Breakout and Seaquest, and obtained a better policy than MuZero on the remaining tasks. Additional results about ReZero-E and EfficientZero are detailed in Appendix C.2, specifically in Figure 6 and 7. For a more comprehensive understanding of our findings, we provide an in-depth comparison of the wall-clock training time up to 100k environment steps for ReZero-M against MuZero in Table 1. The data reveals that, on most games, ReZero require between two to four times less wall-clock time per 100k steps compared to the baseline methods.
5.2 Effect of Reanalyze Frequency
In next two sub-sectons, we will delve into details to explain how ReZero works. Firstly, we will analyze how the Just-in-Time Reanalyze technique manages to simplify the original iterative mini-batch reanalyze scheme to the periodical updated version while maintaining high sample efficiency. Here we adjust the periodical reanalyze frequency(reanalyze the buffer 1 time every $n$ training iterations) in ReZero to a lower or higher value on the MsPacman environment, specifically reanalyze frequency=$\{2,1,1/3,0\}$, where $1$ denotes the setting $n=1000$. The original MuZero variants with the reanalyze ratio of 1 is also included in this ablation experiment as a baseline. Figure 4 shows entire training curves in terms of Wall-time or Env Steps and validates that appropriate reanalyze frequency can save the time overhead without causing any obvious performance loss.
Figure 4: The effect of Reanalyze Frequency in ReZero-M on MsPacman. The horizontal axis in left represents Wall-clock Time (hours) while right side uses Env Steps. The vertical axis indicates Return. Proper reanalyze frequency can improve efficiency while obtaining the comparable Return with MuZero (reanalyze ratio=1).
5.3 Effect of Information Reuse
Indicators Avg. time (ms) tree search (num calls) dynamics model (num calls) data process (num calls)
ReZero-M $\textbf{0.69}\pm\textbf{0.02}$ 6089 122 277
MuZero $1.08\pm 0.09$ $13284$ $256$ $455$
Table 2: Comparisons about detailed indicators between ReZero-M and MuZero inside the tree search. The number of function calls is the cumulative value of 100 training iterations on Pong. Avg. time is the average time of a MCTS across all calls.
To further validate and understand the advantages and significance of the Information Reuse technique proposed in ReZero, we meticulously document a suite of statistical indicators of the tree search process in Table 2. Comparative analysis between ReZero-M and MuZero reveals that the Information Reuse technique reduces the invocation frequency of the dynamics model, the search tree, and other operations like data process transformations. Consequently, this advanced technique in leveraging subsequent time step data contributes to save the tree search time in various MCTS-based algorithms, further leading to overall wall-clock time gains. Besides, Information Reuse also incrementally reallocates computations (tree search simulations) towards previously non-visited nodes rather than repeatedly accessing extant nodes, which maybe yield some performance gains in some environments (e.g. Breakout in Figure 3).
6 Conclusion and Limitation
In this paper, we have delved into the efficiency and scalability of MCTS-based algorithms. Unlike most existing works, we incorporates information reuse and periodic reanalyze techniques to reduces wall-clock time costs while preserving sample efficiency. Our theoretical analysis and experimental results confirm that ReZero efficiently reduces the time cost and maintains or even improves performance across different decision-making domains. However, our current experiments are mainly conducted on the single worker setting, there exists considerable optimization scope to apply our approach into distributed RL training, and our design harbors the potential of better parallel acceleration and more stable convergence in large-scale training tasks. Also, the combination between ReZero and frameworks akin to AlphaZero, or its integration with some recent offline datasets such as RT-X (Padalkar et al., 2023), constitutes a fertile avenue for future research. These explorations could broadening the application horizons of MCTS-based algorithms.
7 Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Proof Materials
To verify the theoretical convergence guarantee of ReZero, we have provided some parts of detailed theoretical proofs in the main paper. In this section, we will present the complete supplementary proofs.
Lemma A.1.
Regret decomposition lemma:Denote $X_{t}$ as the reward at time $t$, $A_{t}$ as the arm chosen at time $t$, and $S_{n}=\sum_{t}X_{t}$, then the regret can be decomposed by regret caused by every sub-optimal arm:
$R_{n}=\sum_{i}E[T_{i}]\Delta_{i}$ (21)
Proof.
Note that
$R_{n}$ $=n\mu^{*}-E[S_{n}]$ (22)
$=\sum_{i}\sum_{t=1}^{n}E[(\mu^{*}-X_{t})\mathbb{I}\{A_{t}=i\}]$ (23)
and
$E[(\mu^{*}-X_{t})\mathbb{I}\{A_{t}=i\}|A_{t}]$ (24)
$=\mathbb{I}\{A_{t}=i\}E[\mu^{*}-X_{t}|A_{t}]$ (25)
$=\mathbb{I}\{A_{t}=i\}(\mu^{*}-\mu_{A_{t}})$ (26)
$=\mathbb{I}\{A_{t}=i\}(\mu^{*}-\mu_{i})$ (27)
$=\mathbb{I}\{A_{t}=i\}\Delta_{i}$ (28)
The proof can be completed by combining the two equations and using the definition of $T_{i}$. ∎
Lemma A.2.
Let $X_{1},...,X_{n}$ be a sequence of independent 1-subgaussian random variables, $\hat{\mu_{t}}=\frac{1}{t}\sum^{t}_{s=1}X_{s}$, $\varepsilon>0$, $a>0$ and
$\kappa=\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{t}+\sqrt{\frac{a}{t^{2}}}\geq\varepsilon\}$ (30)
, then it holds that $E[\kappa]\leq 1+\frac{2\sqrt{a}}{\varepsilon}+\frac{2}{\varepsilon^{2}}$.
Proof.
Take $u$ as $\frac{2\sqrt{a}}{\varepsilon}$, then
$E[\kappa]$ $\leq u+\sum_{t=\lceil u\rceil}^{n}\mathbb{P}(\hat{\mu}_{t}\sqrt{\frac{a}{t^{2}}}\geq\varepsilon)$ (31)
$\leq u+\sum_{t=\lceil u\rceil}^{n}\exp(-\frac{t(\varepsilon-\sqrt{\frac{a}{t^{2}}})^{2}}{2})$ (32)
[…]
$=1+u+\frac{2}{\varepsilon^{2}}=1+\frac{2\sqrt{a}}{\varepsilon}+\frac{2}{\varepsilon^{2}}$ (35)
Proof for Theorem 4.3:
Proof.
Denote $T_{i}(k)$ as the times that arm $i$ has been chosen until time $k$, $\hat{\mu}_{is}$ as the empirical mean based on the first $s$ samples from distribution $P_{i}$, and $\hat{\mu}_{i}(k)=\hat{\mu}_{iT_{i}(k)}$. Then we have:
$T_{i}(n)=\sum_{t=1}^{n}\mathbb{I}\{A_{t}=i\}$ (36)
$\leq\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{1}(t-1)+P_{1}\frac{\sqrt{t-1}}{1+T_{1}(t-1)}\leq\mu_{1}-\varepsilon\}$ (37)
$+\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}$ (38)
for 37, we have
$E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{1}(t-1)+P_{1}\frac{\sqrt{t-1}}{1+T_{1}(t-1)}\leq\mu_{1}-\varepsilon\}]$ (39)
$\leq 1+\sum_{t=2}^{n}\sum_{s=0}^{t-1}\mathbb{P}(\hat{\mu}_{1s}+P_{1}\frac{\sqrt{t-1}}{1+s}\leq\mu_{1}-\varepsilon)$ (40)
$=1+\sum_{t=2}^{n}\sum_{s=1}^{t-1}\mathbb{P}(\hat{\mu}_{1s}+P_{1}\frac{\sqrt{t-1}}{1+s}\leq\mu_{1}-\varepsilon)$ (41)
$\leq 1+\sum_{t=2}^{n}\sum_{s=1}^{t-1}\exp(-\frac{s(\varepsilon+P_{1}\frac{\sqrt{t-1}}{1+s})^{2}}{2})$ (42)
$\leq 1+\sum_{t=2}^{n}\exp(-\frac{1}{2}\varepsilon P_{1}\sqrt{t-1})\sum_{s=1}^{t-1}\exp(-\frac{s\varepsilon^{2}}{2})$ (43)
$\leq 1+\sum_{t=2}^{n}\exp(-\frac{1}{2}\varepsilon P_{1}\sqrt{t-1})\frac{2}{\varepsilon^{2}}$ (44)
$\leq 1+\frac{16}{\varepsilon^{4}P_{1}^{2}}$ (45)
Notes: In 40, since we assume $P_{1}\geq\mu_{1}$, the probability of the term $s=0$ would be 00. Thus, we can discard it. If this assumption doesn’t hold, we can choose to accumulate $t$ starting from a larger $t_{0}$(which satisfies $P_{1}\sqrt{t_{0}-1}>\mu_{1}$ as mentioned in the article). Starting the summation from such a $t_{0}$ ensures all terms of $s=0$ can still be discarded, and all add terms that $t\leq t_{0}$ can be bounded to $1$. This only changes the constant term and won’t affect the growth rate of regret. From 42 to 43, we just need to expand the quadratic term and do some simple inequality scaling. From 43 to 44, we need to notice that $\sum_{s=1}^{t-1}\exp(-\frac{s\varepsilon^{2}}{2})$ is a geometric sequence and scale it to $\frac{2}{\varepsilon^{2}}$. From 44 to 45, we use the inequality $\sum_{t=2}^{n}\frac{1}{e^{a\sqrt{t-1}}}\leq\int_{1}^{\infty}\frac{1}{e^{a\sqrt{t-1}}}$.
for 38, we have
$E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (46)
$\leq E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\sqrt{\frac{n-1}{(1+T_{i}(t-1))^{2}}}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (47)
$\leq 1+E[\sum_{s=1}^{n-1}\mathbb{I}\{\hat{\mu}_{is}+P_{i}\sqrt{\frac{n-1}{(1+s)^{2}}}\geq\mu_{1}-\varepsilon\}]$ (48)
$\leq 1+E[\sum_{s=1}^{n}\mathbb{I}\{\hat{\mu}_{is}-\mu_{i}+P_{i}\sqrt{\frac{n-1}{s^{2}}}\geq\Delta_{i}-\varepsilon\}]$ (49)
with Lemma A.2, we can have
[…]
so
$E[T_{i}(n)]\leq 3+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon}+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}+\frac{16}{\varepsilon^{4}P_{1}^{2}}$ (51)
combined with Lemma A.1, the proof is completed. ∎
Proof for Theorem 4.4:
Proof.
The proof can be obtained by making slight modifications to the proof of 4.3. In case that $\mu_{1}$ is already known,
$T_{i}(n)=\sum_{t=1}^{n}\mathbb{I}\{A_{t}=i\}$ (52)
$\leq\sum_{t=1}^{n}\mathbb{I}\{\mu_{1}\leq\mu_{1}-\varepsilon\}$ (53)
$+\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}$ (54)
$=\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}$ (55)
$\leq 2+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon}+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}$ (56)
In case that a sub-optimal value $\mu_{l}$ is known,
$T_{l}(n)=\sum_{t=1}^{n}\mathbb{I}\{A_{t}=l\}$ (57)
$\leq\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{1}(t-1)+P_{1}\frac{\sqrt{t-1}}{1+T_{1}(t-1)}\leq\mu_{1}-\varepsilon\}$ (58)
$+\sum_{t=1}^{n}\mathbb{I}\{\mu_{l}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}$ (59)
$=\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{1}(t-1)+P_{1}\frac{\sqrt{t-1}}{1+T_{1}(t-1)}\leq\mu_{1}-\varepsilon\}$ (60)
$\leq 1+\frac{16}{\varepsilon^{4}P_{1}^{2}}$ (61)
and the bound of $T_{i}(n)$ for $i\neq l$ keeps unchanged.
In both cases, combined with Lemma A.1, the proof is completed. ∎
Appendix B Implementation Details
B.1 Environments
In this section, we first introduce various types of reinforcement learning environments evaluated in the main paper and their respective characteristics, including different observation/action/reward space and transition functions.
Atari: This category includes sub-environments like Pong, Qbert, Ms.Pacman, Breakout, UpNDown, and Seaquest. In these environments, agents control game characters and perform tasks based on pixel input, such as hitting bricks in Breakout. With their high-dimensional visual input and discrete action space features, Atari environments are widely used to evaluate the capability of reinforcement learning algorithms in handling visual inputs and discrete control problems.
Board Games: This types of environment includes Connect4, Gomoku, where uniquely marked boards and explicitly defined rules for placement, movement, positioning, and attacking are employed to achieve the game’s ultimate objective. These environments feature a variable discrete action space, allowing only one player’s piece per board position. In practice, algorithms utilize action mask to indicate reasonable actions.
B.2 Algorithm Implementation Details
Our algorithm’s implementation is based on the open-source code of LightZero (Niu et al., 2023). Given that our proposed theoretical improvements are applicable to any MCTS-based RL method, we have chosen MuZero and EfficientZero as case studies to investigate the practical improvements in time efficiency achieved by integrating the ReZero boosting techniques: just-in-time reanalyze and speedy reanalyze (temporal information reuse).
To ensure an equitable comparison of wall-clock time, all experimental trials were executed on a fixed single worker hardware setting consisting of a single NVIDIA A100 GPU with 30 CPU cores and 120 GiB memory. Besides, we emphasize that to ensure a fair comparison of time efficiency and sample efficiency, the model architecture and hyper-parameters used in the experiments of Section 5 are essentially consistent with the settings in LightZero (Niu et al., 2023). For specific hyper-parameters of ReZero-M and MuZero on Atari, please refer to the Table 3. The main different hyper-parameters for the ReZero-M algorithm in the Connect4 and Gomoku environment are set out in Tables 4. In addition to employing an LSTM network with a hidden state dimension of 512 to predict the value prefix (Ye et al., 2021), all hyperparameters of ReZero-E are essentially identical to those of ReZero-M in Table 3.
Wall-time statistics Note that all our current tests are conducted in the single-worker case. Therefore, the wall-time reported in Table 1 and Appendix 5 for reaching 100k env steps includes:
  • collect time: The total time spent by an agent interacting with the environment to gather experience data.(Weng et al., 2022) can be integrated to speed up. Besides, this design also makes MCTS-based algorithms compatible to existing RL exploration methods like (Burda et al., 2018).
  • reanalyze time: The time used to reanalyze collected data with the current policy or value function for more accurate learning targets (Schrittwieser et al., 2021).
  • train time: The duration for performing updates to the agent’s policy, value functions and model based on collected data.
  • evaluation time: The period during which the agent’s policy is tested against the environment, separate from training, to assess performance.
Currently, we have set collect_max_episode_steps to 10,000 and eval_max_episode_steps to 20,000 to mitigate the impact of anomalously long evaluation episodes on time. In the future, we will consider conducting offline evaluations to avoid the influence of evaluation time on our measurement of time efficiency. Furthermore, the ReZero methodology represents a pure algorithmic enhancement, eliminating the need for supplementary computational resources or additional overhead. This approach is versatile, enabling seamless integration with single-worker serial execution environments as well as multi-worker asynchronous frameworks. The exploration of ReZero’s extensions and its evaluations in a multi-worker (Mei et al., 2023) paradigm are earmarked for future investigation.
Board games settings Given that our primary objective is to test the proposed techniques for improvements in time efficiency, we consider a simplified version of single-player mode in all the board games. This involves setting up a fixed but powerful expert bot and treating this opponent as an integral part of the environment. Exploration of our proposed techniques in the context of learning through self-play training pipeline is reserved for our future work.
Hyperparameter Value
Replay Ratio (Schwarzer et al., 2023) 0.25
Reanalyze ratio (Schrittwieser et al., 2021) 1
Num of frames stacked 4
Num of frames skip 4
Reward clipping (Mnih et al., 2013) True
Optimizer type Adam (Kingma & Ba, 2015)
Learning rate $3\times 10^{-3}$
Discount factor 0.997
Weight of policy loss 1
Weight of value loss 0.25
Weight of reward loss 1
Weight of policy entropy loss 0
Weight of SSL (self-supervised learning) loss (Ye et al., 2021) 2
Batch size 256
Model update ratio 0.25
Frequency of target network update 100
Weight decay $10^{-4}$
Max gradient norm 10
Length of game segment 400
Replay buffer size (in transitions) 1e6
TD steps 5
Number of unroll steps 5
Use augmentation True
Discrete action encoding type One Hot
Normalization type Batch Normalization (Ioffe & Szegedy, 2015)
Priority exponent coefficient (Schaul et al., 2015) 0.6
Priority correction coefficient 0.4
Dirichlet noise alpha 0.3
Dirichlet noise weight 0.25
Number of simulations in MCTS (sim) 50
Categorical distribution in value and reward modeling True
The scale of supports used in categorical distribution (Pohlen et al., 2018) 300
Table 3: Key hyperparameters of ReZero-M on Atari environments.
Hyperparameter Value
Replay Ratio (Schwarzer et al., 2023) 0.25
Reanalyze ratio (Schrittwieser et al., 2021) 1
Board size 6x7; 6x6
Num of frames stacked 1
Discount factor 1
Weight of SSL (self-supervised learning) loss 0
Length of game segment 18
TD steps 21; 18
Use augmentation False
Number of simulations in MCTS (sim) 50
The scale of supports used in categorical distribution 10
Table 4: Key hyperparameters of ReZero-M on Connect4 and Gomoku environments. If the parameter settings of these two environments are different, they are separated by a semicolon.
Appendix C Additional Experiments
C.1 Sample-efficiency of ReZero-M
Except for the results reported in main paper for time efficiency, we also shows the detailed sample efficiency results in this section. Figure 5 displays the performance over environment interaction steps of the ReZero-M algorithm compared with the original MuZero algorithm across six representative Atari environments and two board games. We can find that ReZero-M obtained similar sample efficiency than MuZero on the most tasks.
Figure 5: Sample-efficiency of ReZero-M vs. MuZero on six representative Atari games and two board games. The horizontal axis represents Env Steps, while the vertical axis indicates the Episode Return over 5 assessed episodes. ReZero achieves similar sample-efficiency than the baseline method. Mean of 5 runs; shaded areas are 95% confidence intervals.
C.2 ReZero-E
The enhancements of ReZero we have proposed are universally applicable to any MCTS-based reinforcement learning approach theoretically. In this section, we integrate ReZero with EfficientZero to obtain the enhanced ReZero-E algorithm. We present the empirical results comparing ReZero-E with the standard EfficientZero across four Atari environments.
Figure 6 shows that ReZero-E is better than EfficientZero in terms of time efficiency. Figure 7 indicates that ReZero-E matches EfficientZero’s sample efficiency across most tasks. Additionally, Table 5 details training times to 100k environment steps, revealing that ReZero-E is 0.5 to 2 times faster than baseline methods on most games.
Figure 6: Time-efficiency of ReZero-E vs. EfficientZero on four representative Atari games. The horizontal axis represents Wall-time (hours), while the vertical axis indicates the Episode Return over 5 assessed episodes. ReZero-E achieves higher time-efficiency than the baseline method. Mean of 3 runs; shaded areas are 95% confidence intervals.
Figure 7: Sample-efficiency of ReZero-E vs. EfficientZero on four representative Atari games. The horizontal axis represents Env Steps, while the vertical axis indicates the Episode Return over 5 assessed episodes. ReZero-E achieves similar sample-efficiency than the baseline method. Mean of 3 runs; shaded areas are 95% confidence intervals.
avg. wall time (h) to 100k env. steps $\downarrow$ Pong MsPacman Seaquest UpNDown
ReZero-E (ours) $\mathbf{2.3}\scriptstyle{\pm 1.4}$ $\mathbf{3}\scriptstyle{\pm 0.3}$ $\mathbf{3.1}\scriptstyle{\pm 0.1}$ $\mathbf{3.6}\scriptstyle{\pm 0.2}$
EfficientZero (Ye et al., 2021) $10\scriptstyle{\pm 0.2}$ $12\scriptstyle{\pm 1.3}$ $15\scriptstyle{\pm 2.3}$ $15\scriptstyle{\pm 0.7}$
Table 5: Average wall-time of ReZero-E vs. EfficientZero on four Atari games.The time represents the average total wall-time to 100k environment steps for each algorithm. Mean and standard deviation over 5 runs. ReZero-E used $\mathbf{0.5-2\times}$ less wall-time per 100k steps than baseline methods on most games.